Two pieces of rope measure 25 feet. One piece is 3 feet more than the other. What are the lengths of the two pieces?
Answer by adamchapman(301)

Call the 2 lengths "x" and "y":
x+y=25
y=x+3..........subtitute this into the first experssion:
x+(x+3)=25
2x+3=25
2x=22
x=11 feet
recalling y=x+5;
y=16 feet
I hope this helps,
Adam
P.S. please visit my website, it may be helpful to you. The address is www.geocities.com/quibowibbler